OpenStudy (anonymous):

A right square pyramid has a slant height of 13 inches and a lateral area of 260 square inches. What is the volume of the pyramid to the nearest cubic inch?

jhonyy9 (jhonyy9):

so like a first step than divide 260 by 4 will get area of a side so than 260/4 = 65 in^2 so you know again that area of a side is base (b)of a triangle time the slant height divide 3 so than will be 65 = b*13/3 so b=65*3/13 b=15| so this mean that the square what is the base of pyramid has sides length equal 15 in you know that the volume of a pyramid is area of base time the height of pyramid divide 3| 

jhonyy9 (jhonyy9):

so area of base will be 15*15 = 225 in^2 the volume of pyramid will be 225*10,61/3 = 795,73 in^3 hope so much that is understandably